A car was going 60 mph and a truck was going 55 mph. They are coming towards each other and are 450 miles apart. They both leave at 5 pm. When will they meet?
Answer by Edwin McCravy(20059)

car was going 60 mph and a truck was going 55 mph. They are coming towards each other and are 450 miles apart. They both leave at 5 pm. When will they meet?
Their approach rate is the sum of their rates = 60+55 = 115 mph

Time = distance/rate = 450/116 = 90/23 = 3 21/23 hours after 5pm

which is 8:54 and 46 22/23 seconds pm or about 8:55 pm.

Edwin